Hello.
I'm studying for a test and need some help checking the following problem:
Solve the equation:
4^x+3 = 8^x-1
thanks!

Answer by stanbon(75887)

Solve the equation:
4^(x+3) = 8^(x-1)
---
[2^2]^(x+3) = [2^3]^(x-1)
---
2^(2x+6) = 2^(3x-3)
---
2x+6 = 3x-1
x = 7
============
Cheers,
Stan H.